Franco Co. reported net sales of $10,000 in year 2 and $8,000 in year 1. Franco reported total assets of $18,000 in year 2 and $22,000 in year 1. Total asset turnover for year 2 would equal (answer should be a decimal number).
Answer #1

Total asset turnover for year 2 is equal to Net Sales divided by the average of total assets.

Net Sales $10,000

Average assets = (18,000+22,000)/2 = 20,000

Total asset turnover = 10,000/20,000 = 0.5